Sodium is a monovalent metal having a density of 0.971 g/cm3 and a molar
mass of 23.0 g/mol. Use this information to calculate (i) the density of charge
carriers and (ii) the Fermi energy of sodium.
1
Expert's answer
2020-12-29T15:32:02-0500

a) ρ = 0.971 g/cm3

M = 23.0 g/mol

"V = \frac{M}{ρ} \\

= \frac{23.0}{0.971} = 23.68 \;cm^3/mol \\

= \frac{23.68}{6.02 \times 10^{23}} = 3.934 \times 10^{-23} \;cm^3"

The density of the charge carriers is

"n_e = \frac{1}{V} \\

= \frac{1}{8.934 \times 10^{-23}} \\

= 2.54 \times 10^{22} \;cm^{-3} \\

= 2.54 \times 10^{28} \;m^{-3}"

b) Fermi energy of sodium is

"E_F = (3.65 \times 10^{-19})n_e^{\frac{2}{3}} \;eV \\

= (3.65 \times 10^{-19})(2.54 \times 10^{28})^{\frac{2}{3}} \;eV \\

= 6.797 \times 0.4645 \;eV \\

= 3.157 \;eV"